Translate this situation into a system of equation of the first degree in two clear variables: To complete a team project, Yvon spent only 3/4 of the time that Peter spent on the project. The difference between the number of hours each of them worked is 20.
Answer by Theo(13342)



let x = the number of hours that yvonne worked.
let y = the number of hours that peter worked.

you get x = 3/4 * y
y - x = 20

those are your two equations that you want to solve simultaneously.

replace x with 3/4 * y in the second equation to get:

y - 3/4 * y = 20

combine like terms to get 1/4 * y = 20

solve for y to get y = 80.

in the equation of y - x = 20, replace y with 80 to get 80 - x = 20 and solve for x to get x = 60.

you get y = 80 and x = 60

y is the number of hours that peter worked.
x is the number of hours that yvonne worked.

x = 3/4 * y
y - x = 20

the requirements of the problem were satisfied and that's your solution.